OpenStudy (anonymous):

I need to know how to solve the equation in the interval [0, 2 ]. cos(t)= -sq rt of 3/2

OpenStudy (anonymous):

\[\cos t=-\frac{ \sqrt{3} }{ 2 }=-\cos \frac{ \pi }{ 6 }\] now you know where cost is negative.